OpenStudy (anonymous):

ln 1/e^3 I got -1 but Im not sure if that is correct?

OpenStudy (jtvatsim):

almost, I think you are on the right track though... let me work it out first... :)

OpenStudy (anonymous):

okay

OpenStudy (jtvatsim):

alright, so first convert 1/e^3 to negative exponents, what did you get (or does that even make sense :) )?

OpenStudy (anonymous):

thats where I got the -1, bc if its a fraction it has to come from a neg number right?

OpenStudy (jtvatsim):

or wait, that exponent might be different that thinking...

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Charlie how did you get the answer?

OpenStudy (anonymous):

ln(e^-3), ln and e cancels each other

OpenStudy (jtvatsim):

is the question this or this? \[\ln(\frac{1}{e^3})\] \[[\ln(1/e)]^3\]

OpenStudy (jtvatsim):

right, I think charlie's got it :)

OpenStudy (anonymous):

the first

OpenStudy (jtvatsim):

ok, then yes, the conversion should be e^-3 because not only is it a fraction (-1) but it is also to the third power (3) so combining gives you -3

OpenStudy (anonymous):

so when I write it out I can write it like this?

OpenStudy (anonymous):

Yes, both ways

OpenStudy (anonymous):